OpenStudy (anonymous):

A small object with mass 3.90 kg moves counterclockwise with constant speed 5.00 m/s in a circle of radius 3.90 m centered at the origin. It starts at the point with position vector (3.90 + 0 ) m. Then it undergoes an angular displacement of 9.00 rad. (a) What it its position vector? ( + ) m (b) In what quadrant is the particle located and what angle does its position vector make with the positive x-axis? at ° (c) What is its velocity? ( Your response differs from the correct answer by more than 10%. Double check your calculations. + ) m/s (d) In what d

OpenStudy (anonymous):

hi so its moving negative so thus makes it positive, with me?

OpenStudy (anonymous):

have you drawn a free hand diagram yet?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how many radians are in a circle

OpenStudy (anonymous):

lets have you draw one on screen you draw one and then i'll draw one

OpenStudy (anonymous):

the question is a stop watch has a circular dail divided into 120 division.what is its least count

OpenStudy (anonymous):

so a cirlce is 2pi radians so 360=2pi radians, ok

OpenStudy (anonymous):

do you know the angular accelration formulal

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you need to calculate the distance traveled or x or s depending on your formula sheet that the 3.9kg travels when traveling at 5.0 m/s

OpenStudy (anonymous):

here only i have to know the formula for this

OpenStudy (anonymous):

here is a formula you can try. \[\alpha=T/mr ^{2}\]

OpenStudy (anonymous):

T is not time, but its Torque

OpenStudy (anonymous):

example

OpenStudy (anonymous):

so you need find your force. F=ma

OpenStudy (anonymous):

so your mass 3.90kg x 9.8m/s

OpenStudy (anonymous):

38.22 N

OpenStudy (anonymous):

take F and multily it by radius, that gives you T

OpenStudy (anonymous):

so T/mr^2=alpha or angular accelration

OpenStudy (anonymous):

T=152.96

OpenStudy (anonymous):

ok i understand thanks

OpenStudy (anonymous):

so alpha=152.96/3.9(3.9^2)

OpenStudy (anonymous):

good from there you should be able to find the other components.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome